assuming that y varies directly as x, if y = 50 when x = 8, find x when y = 75
Answer by MsRedz123(24)

If y varies directly as x. Then, y=kx where k is constant.
y=kx
50=k(8)
50/8=k(8)/8
k=50/8 or 25/4
y=kx
75=(25/4)(x)
75/(25/4)=x
x=75(4/25)
x=300/25
x=12
Therefore, x is 12.